OpenStudy (anonymous):

6( x + 8) -6(-3 + 1 y) = -5. solve for slope and x and y intercept please explain through showing your work, I've kind of hit a wall on this one.

OpenStudy (anonymous):

6x+48+18-6y=-5 6x+66+5=6y 6x+71=6y y=6/6x+71/6 y=x+71/6 which may be compared to y=mx+c where m is the slope of line here,m=1

OpenStudy (anonymous):

6( x + 8) -6(-3 + 1 y) = -5. i.e. 6 x + 48 +18 - 6 y = -5. i.e. 6 x -6y + 66 = -5. i.e. 6 x -6y = -5-66 i.e. 6 x -6y = -71 i.e. 6 x +71 = 6y i.e 6y = 6x+71 \[y= x+\frac{71}{6}\] comparing it with y=mx+b we find slope m= 1 and y-intercept b= 71/6 for x-intercept, substitute y=0 in the eq we find \[y= x+\frac{71}{6} \rightarrow 0 = x+\frac{71}{6} \rightarrow x= -\frac{71}{6}\] Hence slope m= 1 x-intercept = -71/6 y-intercept = 71/6 @JustinSant

OpenStudy (anonymous):

Thanks guys, this helped a lot I just had the negative symbols switched.